OpenStudy (anonymous):

Write an equation of a horizontal line passing through the two points (-1,5) and (0,5). Do not use any spaces in your answer.

OpenStudy (aaronq):

A horizontal line has a slope of zero, and looks like this If you need to show some work, use: \(\sf (y_1-y)=m(x_1-x)\) Pick a point (-1,5); m=slope =0 \(\sf (5-y)=0(-1-x)\) \(\sf 5-y=0 \) \(\sf y=5 \)

OpenStudy (anonymous):

Thank-you so much :)) @aaronq